Small pieces were three dollars and large pizzas for five dollars. With the throne, it was necessary to spend $475 for 125 pizzas. How many pieces of each type of purchased?
1
Expert's answer
2019-05-20T03:21:20-0400

Assume they bought "x" small pizzas for $3 and "y" large pizzas for $5. Then we can write a system of equations:


"x+y=125,""3x+5y=475."

Solve by subtraction. Multiply both sides of the first equation by 5:


"5x+5y=125\cdot5.""3x+5y=475."

Subtract the left side of the second from the left side of first, do this for the right sides:


"5x+5y-3x-5y=625-475."

We've got simple equation for "x":


"2x=150,\\
x=75."


Substitute this to the first one and find "y":


"y=125-75=50."

75 small pizzas and 50 large ones.